### v1

# Tuning Colloidal Reactions

Abstract
The precise control of complex reactions is critical for biological processes ranging from cell division to metabolism. Synthetic analogues of living materials suffer from our inability to tune chemical reactions with precise outcomes. Here, we leverage differentiable simulators to design nontrivial reaction pathways in colloidal assemblies. By optimizing interactions between reactants and substrates, we achieve controlled disassembly of octahedral and icosahedral shells. As a potential engineering target, we design a reaction that provokes the release of a small particle trapped in a shell.
Both living and non-living physical systems exhibit complex, dynamical behavior, ranging from repair, locomotion, and catalysis. Fundamentally, these complex dynamical behaviors arise from sequences of reactions, in which a set of substances (i.e. the reactants) are transformed into a set of different substances (i.e. the products).
A rich body of theoretical work aims to characterize and tune systems of interacting agents spanning a range of system descriptions [1, 2, 3]. There have been exciting advances in creating tunable reactions that are experimentally realizable [4, 5, 6, 7, 8, 9]. However, the models that provide an experimentally-relevant level of detail are typically bespoke and application-specific, necessitating entirely new research programs for each new reaction. Conversely, general theoretical models have provided deep insights into the basis of self-limiting growth [10, 11], self-replication [12],catalysis [13, 14], and much more [15, 16, 17, 18, 19]. These models, however, are largely too abstract to inform experimental design. Moreover, many reactions that yield desirable dynamical behavior are incredibly finely-tuned, making engineering these reactions untenable with most traditional design approaches. To design precise reactions, it is necessary to carry out inverse design, whereby one designs components and their interactions to achieve a given reaction. While inverse design has successfully applied to self assembly [20, 21, 22, 23, 24], its application to reaction pathways is much more difficult owing to the need to choose design parameters to favor particular dynamical trajectories. The advent of differentiable simulators [24], powered by software libraries developed for machine learning [25], has opened up the possibilty of directly designing reactions.
Here, we use differentiable molecular dynamics (MD) to carry out the inverse design of complex reactions. The differentiability makes it possible to compute analytical gradients with respect to trajectories during the interaction of components and perform gradient-based optimization with respect to parameters to achieve a target reaction. To demonstrate the ability framework to design nontrivial reactions, we consider the controlled disassembly of colloidal structures, whereby one particle is extracated from an otherwise complete shell of colloidal particles. Disassembly is central to the dynamic functions of living systems, such as defect repair, self-replication, and catalysis. Most existing examples of controlled disassembly in synthetic systems rely on external forcing to drive the disassembly process [26, 27, 28, 29], which provides a direct pathway to tuning behavior. However, for many engineering applications including those inherent to living systems, the use of external fields is limiting. Alternatively, living systems typically rely on local energy consumption (e.g. biological enzymes consuming ATP) rather than global fields, but the synthetic design of these systems is significantly more complex.
In this paper, we design for the controlled disassembly of both an octahedral shell and an icosahedral shell. Importantly, our disassembly mechanism does not rely on external forcing and occurs without consuming energy. As a model for potential engineering applications, we apply our disassembly mechanism to release a small particle trapped inside an icosahedral shell. Controlled disassembly serves as a striking example of a complex reaction because the reaction requires a finely-tuned interaction energy to keep the remaining shell stabilized while performing the desired disassembly. Moreover, we consider a fixed substrate and only parameterize an external structure that acts on this fixed substrate, enabling control over disassembly without modifying critical components of the target reaction.
Results.—We implement controlled disassembly example in a patchy particle system. Patchy particles have long been used to emulate interactions in soft materials[30, 31], and offer tremendous flexibility of designed interactions, especially when patches can be programmed to interact with each other with designed strength and range. Optimizing patchy particle systems to achieve specific behaviors has been made possible by the recent development of patchy particle simulations within a differentiable library [24].The ability to carry out this type of optimization within a molecular dynamics simulations opens up enormous possibilities for the design of novel reactions. In particular, we aim to remove a single particle from a shell composed of patchy particles without disrupting the remaining shell structure. We tune disassembly without changing any properties of the shell itself, by introducing an external reactant that interacts with the shell to disassemble it in the desired manner. We term the external reactant a “spider” due to the structure of the object.
The shell is modelled as a collection of patchy particles where each patch corresponds to a contact with a neighboring particle. Patches interact via a Morse potential (, ) and vertex centers interact via soft-sphere repulsion (). Importantly, the geometry and interaction energy of a shell vertex are fixed throughout an optimization: only features of the spider are optimized.
The spider is modeled as a rigid body composed of a ring of “base” particles and a “head” particle that sits above the ring along its symmetry axis. The base particles are connected by repulsive bars and the head is connected to the base particles particles by the same repulsive bars, making the entire structure a cage-like object that is open on one end. The head particle interacts with shell vertices via a Morse potential whereas base particles and connecting bars interact with shell vertices via soft sphere repulsion. Unlike the shell, the geometry and interaction energy of the spider are parameters of the optimization. See Figure 1C for an overview of this parameterization. All the interaction energies in our system are parameterized with simple, physics-based potentials.
We optimize over 8 parameters that characterize the geometry of the spider and its interaction with the shell via differentiable MD (see Figure 1D). Our loss term is constructed via two competing terms: one that rewards a final state in which the target vertex is extracted, and one that penalizes a strong interaction between the spider and non-target vertices. Consider a shell comprised of a collection of vertices where and for the octahedron and icosahedron, respectively. We seek to extract a target vertex from the shell while leaving the remaining shell intact. We can measure the degree to which is successfully extracted via the following expression:
where denotes the Euclidean distance between vertices and . Note that the negative sign as we formulate our optimization problem to minimize the loss. Secondly, we can minimize the interaction energy between the spider and the non-target vertices of the shell,
where represents the interaction energy between the spider head, , and a shell vertex as depicted in Figure 1C. For the total loss, we evaluate the interaction between the spider and the non-target shell vertices with respect to the initial state while we calculate the spider’s extraction of the target vertex with respect to the final state. We evaluate the remaining energy term on the initial state because this term is only well defined when evaluated on the same configuration at each iteration. In all simulations, the spider is initialized bound to the target vertex and the system is integrated for 1000 timesteps (see Appendix).
Octahedron.—We first demonstrate our method on an octahedron, a simple but nontrivial platonic solid. We explore two limits of our optimization procedure: weak and strong initial shell-spider interactions. Figure 2 summarizes our results.
First, we perform an optimization where the spider is initialized to interact weakly with the shell vertices (, ). Under this weak initial interaction, the spider simply diffuses away from the shell at long timescales without extracting the target vertex. Then, throughout the course of the optimization, we observe variable changes consistent with increasing the interaction between the spider head and the shell – increases, the head height decreases, and the head particle radius increases. However, we also observe less intuitive changes – increases (decreasing the range of the Morse potential), the shell base radius decreases, and the base particle radius increases. This suggests tightly coupled, nontrivial parameter changes drive the extraction of the target vertex while maintaining minimal interaction with the remaining shell.
Next, we performed an optimization in the opposite limit in which the spider is initialized to interact strongly with the shell (, ). Initially, this interaction is so strong that the spider not only extracts the target vertex but it also disrupts the remaining shell. This disruption can be seen in the large value of the remaining energy loss term, which penalizes the energy between the spider head and non-target vertices. Throughout the optimization, we observe variable changes consistent with tuning the interaction strength to maintain extraction while minimizing off-target interactions – decreases, increases, the head radius decreases, the head height increases, and the base particle radius increases. When evaluated on longer simulations, the converged parameter set also achieves diffusion of the spider-vertex complex from the remaining, undisturbed shell.
While only single optimizations are presented in Figures 2 and 3, we additionally performed ensembles of optimizations over different initial seeds for the random number generator and over different random perturbations to the initial parameter in both the high and low energy regimes. Despite the wide variance in initial parameters, we observe low variance in the converged parameter values in all cases (see Figure 4). This may suggest that the solution set is narrow.
Contrasting these two optimization regimes reveals the inherent delicacy in tuning the spider to achieve extraction of a target vertex and subsequent diffusion of the vertex away from the shell. The spider-shell interaction must be sufficiently strong to extract the target vertex, but be simultaneously weak enough to not disturb the non-target vertices and to diffuse away from the shell within the timescale of our simulations. This tension can be seen in the behavior of the loss terms in each optimization. In the weak-interaction limit, the term penalizing interactions with non-target vertices remains negligible while the extraction term drives optimization; in the strong-interaction limit it is the same energy-penalizing term that dominates the loss.
Icosahedron and Controlled Release.—We now increase in complexity to a model of a biologically relevant function: controlled release of a particle from an icosahedral shell. We initialize our system with a small particle trapped inside a fully formed icosahedral shell. We then optimize a similar spider structure to perform controlled disassembly of the icosahedral shell, resulting in the controlled release of the small particle. This serves as a toy example of a potential target for engineering applications, such as drug delivery via a viral shell.
To match the symmetry of the icosahedron, we use a spider with five legs rather than four. Otherwise, optimization proceeds in the same fashion as for the octahedral shell. We consider the same two limits: one with high interaction strength between the spider and the shell and one with low interaction strength between the spider and the shell (see Figure 3). The same delicate tradeoff between extraction and diffusion manifests with the icosahedron, and our automatic differentiation-based optimization deftly navigates this tradeoff.
Discussion.—In this Letter, we have demonstrated the design of nontrivial reactions via differentiable simulation. We considered the case of controlled disassembly, in which there is an inherent tension between initiating disassembly and maintaining the integrity of the remaining substructure. We first consider the controlled disassembly of an octahedral shell composed of patchy particles and show how the parameters governing an external structure (i.e. the “spider”) can be finely tuned to minimize a loss function representing this inherent tension. We then increase in complexity, demonstrating the controlled release of a small particle from an icosahedral shell.
By optimizing directly with respect to the numerically integrated dynamics, our method is general enough to study a wide range of systems. Foremost, our method may enable the experimental realization of theoretical models that were otherwise limited by an inability to finely tune interaction energies. For example, Ref. [12] introduces a model of self-replicating colloidal clusters in which kinetic traps can be avoided by tuning the interaction energies, but dissociation of a new cluster from its parent (a necessary step for replication) required an artificial trigger event in numerical simulations. In contrast, our designed parameters lead to spontaneous dissociation of the spider-vertex complex away from the remaining shell. The computational flexibility of the method could also easily allow restricting the parameter regime to experimentally realizable interactions. This could be done for DNA coated colloids, for instance, by optimizing the DNA sequences that define the interaction strength [32, 33].
A number of numerical instabilities can arise while optimizing for the types of reactions we consider. The primary limitation we observe is that gradients become unstable and very large for long simulations. There are several possible approaches to reducing instability in gradients of long simulations. One standard method to mitigate such instabilities in the context of differentiable programming is gradient clipping [34, 35]. One could also decrease the total number of timesteps by training an emulator to resolve the dynamics with a larger timestep than is otherwise possible with standard integrators, following similar work for deterministic systems [36, 37, 38]. An alternative approach to decreasing the number of timesteps would be to integrate differentiable simulations with enhanced sampling methods. As a result, one could sample low probability events without the need for long simulation times.
We rely on gradient-based optimization due to its scalability and performance. The method scales naturally to larger systems since the calculation of a gradient via automatic differentiation only requires a single simulation, while at the same time reverse-mode automatic differentiation scales efficiently with the number of parameters [39]. Calculating the analytical gradient with respect to a simulation explicitly captures interdependency between parameters, essential to efficiently tuning complex behavior.
The experimental realization of theoretical models of complex dynamical behavior has long been limited by the ability to finely tune the parameterization of such models. This method serves as a potential bridge between these two worlds, permitting the delicate tradeoffs inherent in parameterizations of complex reactions to be navigated efficiently using physics-based potentials and dynamics.
Acknowledgements.
We thank Sam Schoenholz for his work developing JAX-MD and the members of the Brenner Group at Harvard University for helpful discussions. This material is based upon work supported by the Office of Naval Research (ONR N00014-17-1-3029, ONR N00014-23-1-2654), the NSF Grant DMR-1921619, the NSF AI Institute of Dynamic Systems (#2112085), and a grant from the Simons Foundation (#1141499).References
- van Anders et al. [2015] G. van Anders, D. Klotsa, A. S. Karas, P. M. Dodd, and S. C. Glotzer, Digital alchemy for materials design: Colloids and beyond, Acs Nano 9, 9542 (2015).
- Zeravcic et al. [2017] Z. Zeravcic, V. N. Manoharan, and M. P. Brenner, Colloquium: Toward living matter with colloidal particles, Reviews of Modern Physics 89, 031001 (2017).
- Sherman et al. [2020] Z. M. Sherman, M. P. Howard, B. A. Lindquist, R. B. Jadrich, and T. M. Truskett, Inverse methods for design of soft materials, The Journal of chemical physics 152 (2020).
- Truong et al. [2022] V. X. Truong, L. L. Rodrigues, and C. Barner-Kowollik, Light-and mechanic field controlled dynamic soft matter materials, Polymer Chemistry 13, 4915 (2022).
- Li et al. [2019] Q. Li, A. P. Schenning, and T. J. Bunning, Light-responsive smart soft matter technologies (2019).
- Wang et al. [2012] Y. Wang, Y. Wang, D. R. Breed, V. N. Manoharan, L. Feng, A. D. Hollingsworth, M. Weck, and D. J. Pine, Colloids with valence and specific directional bonding, Nature 491, 51 (2012).
- McMullen et al. [2022] A. McMullen, M. Muñoz Basagoiti, Z. Zeravcic, and J. Brujic, Self-assembly of emulsion droplets through programmable folding, Nature 610, 502 (2022).
- Niu et al. [2019] R. Niu, C. X. Du, E. Esposito, J. Ng, M. P. Brenner, P. L. McEuen, and I. Cohen, Magnetic handshake materials as a scale-invariant platform for programmed self-assembly, Proceedings of the National Academy of Sciences 116, 24402 (2019).
- Schade et al. [2013] N. B. Schade, M. C. Holmes-Cerfon, E. R. Chen, D. Aronzon, J. W. Collins, J. A. Fan, F. Capasso, and V. N. Manoharan, Tetrahedral colloidal clusters from random parking of bidisperse spheres, Physical review letters 110, 148303 (2013).
- Murugan et al. [2015a] A. Murugan, J. Zou, and M. P. Brenner, Undesired usage and the robust self-assembly of heterogeneous structures, Nature communications 6, 6203 (2015a).
- Zeravcic et al. [2014] Z. Zeravcic, V. N. Manoharan, and M. P. Brenner, Size limits of self-assembled colloidal structures made using specific interactions, Proceedings of the National Academy of Sciences 111, 15918 (2014).
- Zeravcic and Brenner [2014] Z. Zeravcic and M. P. Brenner, Self-replicating colloidal clusters, Proceedings of the National Academy of Sciences 111, 1748 (2014).
- Zeravcic and Brenner [2017] Z. Zeravcic and M. P. Brenner, Spontaneous emergence of catalytic cycles with colloidal spheres, Proceedings of the National Academy of Sciences 114, 4342 (2017).
- Muñoz-Basagoiti et al. [2023] M. Muñoz-Basagoiti, O. Rivoire, and Z. Zeravcic, Computational design of a minimal catalyst using colloidal particles with programmable interactions, Soft Matter 19, 3933 (2023).
- Keim et al. [2019] N. C. Keim, J. D. Paulsen, Z. Zeravcic, S. Sastry, and S. R. Nagel, Memory formation in matter, Reviews of Modern Physics 91, 035002 (2019).
- Murugan et al. [2015b] A. Murugan, Z. Zeravcic, M. P. Brenner, and S. Leibler, Multifarious assembly mixtures: Systems allowing retrieval of diverse stored structures, Proceedings of the National Academy of Sciences 112, 54 (2015b).
- Holmes-Cerfon et al. [2013] M. Holmes-Cerfon, S. J. Gortler, and M. P. Brenner, A geometrical approach to computing free-energy landscapes from short-ranged potentials, Proceedings of the National Academy of Sciences 110, E5 (2013).
- Holmes-Cerfon [2016] M. C. Holmes-Cerfon, Enumerating rigid sphere packings, siam REVIEW 58, 229 (2016).
- Murugan and Vaikuntanathan [2017] A. Murugan and S. Vaikuntanathan, Topologically protected modes in non-equilibrium stochastic systems, Nature communications 8, 13881 (2017).
- Torquato [2009] S. Torquato, Inverse optimization techniques for targeted self-assembly, Soft Matter 5, 1157 (2009).
- Chen et al. [2018] D. Chen, G. Zhang, and S. Torquato, Inverse design of colloidal crystals via optimized patchy interactions, The Journal of Physical Chemistry B 122, 8462 (2018).
- Rechtsman et al. [2006] M. Rechtsman, F. Stillinger, and S. Torquato, Designed interaction potentials via inverse methods for self-assembly, Physical Review E 73, 011406 (2006).
- Ma and Ferguson [2019] Y. Ma and A. L. Ferguson, Inverse design of self-assembling colloidal crystals with omnidirectional photonic bandgaps, Soft matter 15, 8808 (2019).
- King et al. [ress] E. M. King, C. X. Du, Q.-Z. Zhu, S. S. Schoenholz, and M. P. Brenner, Programmable patchy particles for materials design, Proceedings of the National Academy of Sciences (in press).
- Bradbury et al. [2018] J. Bradbury, R. Frostig, P. Hawkins, M. J. Johnson, C. Leary, D. Maclaurin, G. Necula, A. Paszke, J. VanderPlas, S. Wanderman-Milne, and Q. Zhang, JAX: composable transformations of Python+NumPy programs (2018).
- Jung et al. [2023] Y.-j. Jung, H. Kim, H.-K. Cheong, and Y.-b. Lim, Magnetic control of self-assembly and disassembly in organic materials, Nature Communications 14, 3081 (2023).
- Tottori et al. [2013] S. Tottori, L. Zhang, K. E. Peyer, and B. J. Nelson, Assembly, disassembly, and anomalous propulsion of microscopic helices, Nano letters 13, 4263 (2013).
- Martínez-Pedrero et al. [2020] F. Martínez-Pedrero, F. Ortega, J. Codina, C. Calero, and R. G. Rubio, Controlled disassembly of colloidal aggregates confined at fluid interfaces using magnetic dipolar interactions, Journal of colloid and interface science 560, 388 (2020).
- Kostiainen et al. [2011] M. A. Kostiainen, P. Ceci, M. Fornara, P. Hiekkataipale, O. Kasyutich, R. J. Nolte, J. J. Cornelissen, R. D. Desautels, and J. van Lierop, Hierarchical self-assembly and optical disassembly for controlled switching of magnetoferritin nanoparticle magnetism, Acs Nano 5, 6394 (2011).
- Zhang and Glotzer [2004] Z. Zhang and S. C. Glotzer, Self-assembly of patchy particles, Nano letters 4, 1407 (2004).
- Glotzer and Solomon [2007] S. C. Glotzer and M. J. Solomon, Anisotropy of building blocks and their assembly into complex structures, Nature materials 6, 557 (2007).
- Rogers [2020] W. B. Rogers, A mean-field model of linker-mediated colloidal interactions, The Journal of chemical physics 153 (2020).
- Rogers and Crocker [2011] W. B. Rogers and J. C. Crocker, Direct measurements of dna-mediated colloidal interactions and their quantitative modeling, Proceedings of the National Academy of Sciences 108, 15687 (2011).
- Greener [2023] J. G. Greener, Differentiable simulation to develop molecular dynamics force fields for disordered proteins, bioRxiv , 2023 (2023).
- Engel et al. [2022] M. C. Engel, J. A. Smith, and M. P. Brenner, Optimal control of nonequilibrium systems through automatic differentiation, arXiv preprint arXiv:2201.00098 (2022).
- Sanchez-Gonzalez et al. [2020] A. Sanchez-Gonzalez, J. Godwin, T. Pfaff, R. Ying, J. Leskovec, and P. Battaglia, Learning to simulate complex physics with graph networks, in International conference on machine learning (PMLR, 2020) pp. 8459–8468.
- Allen et al. [2022] K. R. Allen, T. Lopez-Guevara, K. Stachenfeld, A. Sanchez-Gonzalez, P. Battaglia, J. Hamrick, and T. Pfaff, Physical design using differentiable learned simulators, arXiv preprint arXiv:2202.00728 (2022).
- Bar-Sinai et al. [2019] Y. Bar-Sinai, S. Hoyer, J. Hickey, and M. P. Brenner, Learning data-driven discretizations for partial differential equations, Proceedings of the National Academy of Sciences 116, 15344 (2019).
- Baydin et al. [2018] A. G. Baydin, B. A. Pearlmutter, A. A. Radul, and J. M. Siskind, Automatic differentiation in machine learning: a survey, Journal of Marchine Learning Research 18, 1 (2018).